| Name | Email | Date | Comment
|
| ANS
| anolla@naic.edu
|
11/4/99
| There it was a bad PIN diode on the first T/R sw causing chn. B not to be seen. VI
checked it.
|
| ANS
| anolla@naic.edu
|
7/26/2001
| The recvrs were hooked to RA , they were installed for radar mode.
|
| ANS
| anolla@naic.edu
|
11/9/2001
| There is not 47 MHz receiver due that the L.O. synth was removed in order to use
it for a test with the 430 MHz receiver at the Dome.
|
| ANS
| anolla@naic.edu
|
1/21/2002
| L.O. synth back into system , but is off due to RFI caused by it.In case that
somebody want to use it we have to go up and turn the synth on.
|
| ANS
| anolla@naic.edu
|
4/3/2002
| The receiver LO frequency is 76.8 Mhz .
|
| ans | anolla@naic.edu | Friday, August 9, 2002 08:38:31 | The L.O. Synth is off to avoid RFI.
|
| ans | anolla@naic edu. | Monday, March 10, 2003 13:12:02 | The L.O. synth was on since January.It was turned off.
|
| ans | anolla@naic.edu | Wednesday, August 11, 2004 15:06:58 | Receiver set for xmitter mode. L.O. synth. off , it causes rfi to the lb receivers.
|
| ans | anolla@naic.edu | Monday, August 16, 2004 09:57:56 | System set up for a radar run , Synth. on.
|
| ans | anolla@naic.edu | Friday, October 8, 2004 13:29:13 | The 47 mhz receiver had the L.O. synth off ,was turned on.
|
| ans | anolla@naic.edu | Thursday, October 21, 2004 14:16:50 | 47 mhz receiver L.O. on for T1998 run tomorrow.
|
| ans | anolla@naic.edu | Wednesday, November 17, 2004 10:54:18 | L.O. synth is off.
|
| ans | anolla@naic.edu | Thursday, December 30, 2004 14:27:04 | System set for R.A. . Synth on.
|
| ans | anolla@naic.edu | Tuesday, January 25, 2005 13:05:03 | L.O. synth is off in order to avoid rfi coming from the 47 Mhz recvr.It causes rfi at the LBW recvr.
|
| ans | anolla@naic.edu | Tuesday, February 8, 2005 15:49:35 | The 47 Mhz L.O. synth is on.
|
| ans | anolla@naic.edu | Wednesday, April 27, 2005 14:14:02 | System set to R.A. mode, synth. ON. System under operational R.A. conditions.
|
| BJ | wgenter@naic.edu | Friday, May 6, 2005 10:36:17 | LO off until 15 May 2005.
|
| ans | anolla@naic.edu | Wednesday, May 18, 2005 09:01:33 | System is off , Dan cancelled its run with it.
|
| W. Genter | wgenter@naic.edu | Monday, July 18, 2005 14:10:39 | Found CH B non operational this AM during System checks. Changed CH A IF output for CH B and found that there was no output in CH A. Problem is in the cable or IF AMP in Receiver Room.
|
| W. Genter | wgenter@naic.edu | Monday, July 18, 2005 14:56:12 | The problem with CH B located and repaired. Evidently the CH B cable had been to removed to work on the chassis (paint) and the CH B cable had not been securely attached.
|
| ans | anolla@naic.edu | Monday, November 21, 2005 11:16:30 | The recvr was checked , it is in the R.A. mode.The synth L.O. is off to avoid RFI.
|
| ans | anolla@naic.edu | Thursday, January 19, 2006 13:59:11 | On 1/17/2006 WG found that the receiver was dead.It was found that somebody removed the HP LO synth ac power from the UPS , so there it was a power dip that shut off the synth rf power for not been in the UPS.Recvr checked this morning and was ok.
|
| ans | anolla@naic.edu | Thursday, January 19, 2006 14:03:18 | Forgot to mention that the recvrs cal are the followings:
Channel A cal value is 5700 k deg & channel B cal value is 4792 k deg. Tsys for both channels normaly =/- 800 k deg. A/L change is =/- 10 db.
|
| ans | anolla@naic.edu | Tuesday, January 24, 2006 15:33:16 | There was a complaint that the system was not phase locking.JRV found that the X2 multiplier P.S. for the reference signal of the HP synth for the LO was bad.
|
| ans | anolla@naic.edu | Thursday, February 2, 2006 09:40:24 | Switching power supply replaced at the X 2 reference multiplier for the LO synt.
|
| Denis Urbain | durbain@naic.edu | Thursday, November 15, 2012 13:02:09 | Problem with the 76.8Mhz reference generator, the RF output switch is not working well. Even with the RF switched ON sometimes the output stays OFF. You have to cycle the RF output switch several times to make it work.
The generator will have to be replaced or send back for repair.
|